OpenStudy (anonymous):

what are the roots to 3x^2-27x-0 ?!

OpenStudy (anonymous):

you could use the quadratic formula to find the roots : i got the formula on a attachment

OpenStudy (accessdenied):

Also, you could set it equal to 0 and solve for x by factoring (usually factoring is easier than the formula, when possible): 3x^2 - 27x - 0 = 0 - 0 just disappears... 3x^2 - 27x = 0 Greatest Common Factor: 3x 3x(x - 9) = 0 If ab = 0, then a = 0 or b = 0 3x = 0 or x - 9 = 0 You can solve for x in each case for the roots.

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Access Denied provided the formulas: 3x = 0 and x-9 = 0 Solving for x: 3x/3 = 0/3 , x = 0 x - 9 + 9 = 0 + 9, x = 9 The solution indicates that the parabola crosses the x axis at 0 and also at +9.